OpenStudy (anonymous):

Use an addition property to solve for c. 14 + (–12) = –12 + c –26 –14 14 26

OpenStudy (alexandervonhumboldt2):

14+(-12)=14-12=?? can you irst solve this one?

OpenStudy (anonymous):

simplify the left side first

OpenStudy (smilie2015):

14+(-2)=-12+c 2+12=c 14=c

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Add twelve to both sides 2+12=?

OpenStudy (anonymous):

x=? You should get this by adding twelve to both sides

OpenStudy (anonymous):

x=14